# Quantifying Calibration Error in Modern Neural Networks through Evidence Based Theory

### v1

@@ Abstract
Trustworthiness in neural networks is crucial for their deployment in critical applications, where reliability, confidence, and uncertainty play pivotal roles in decision-making. Traditional performance metrics such as accuracy and precision fail to capture these aspects, particularly in cases where models exhibit overconfidence. To address these limitations, this paper introduces a novel framework for quantifying the trustworthiness of neural networks by incorporating subjective logic into the evaluation of Expected Calibration Error (ECE). This method provides a comprehensive measure of trust, disbelief, and uncertainty by clustering predicted probabilities and fusing opinions using appropriate fusion operators. We demonstrate the effectiveness of this approach through experiments on MNIST and CIFAR-10 datasets, where post-calibration results indicate improved trustworthiness. The proposed framework offers a more interpretable and nuanced assessment of AI models, with potential applications in sensitive domains such as healthcare and autonomous systems.
Keywords AI Trustworthiness  $\cdot$ Neural networks  $\cdot$ Calibration  $\cdot$ Expected Calibration Error (ECE)  $\cdot$ Subjective Logic  $\cdot$ Uncertainty Quantification  $\cdot$ Temperature Scaling  $\cdot$ MNIST  $\cdot$ CIFAR-10  $\cdot$ AI Reliability.
1 Introduction
Artificial Intelligence (AI) systems, particularly neural networks, are increasingly employed in critical applications such as healthcare, finance, and autonomous systems. These systems play an integral role in decision-making, where the trustworthiness of their predictions becomes paramount. Trustworthiness in AI encompasses attributes like reliability, robustness, fairness, and transparency, yet these qualities are often difficult to evaluate, particularly in neural networks, which are typically viewed as "black-box" models. This opacity raises significant concerns about their trustworthiness, especially in sensitive domains where incorrect decisions can lead to severe consequences.
@@ 1 Introduction
Problem. Evaluating the trustworthiness of an AI system, particularly neural networks, poses significant challenges. Traditional metrics like accuracy, precision, and recall do not capture the uncertainties and subjective aspects of trustworthiness. Confidence calibration methods, such as those discussed in Guo et al. (2017), introduce the Expected Calibration Error (ECE) to align predicted probabilities with actual outcomes, aiming to make AI systems more trustworthy. However, ECE, while useful, is not easily interpretable and does not provide a comprehensive view of an AI system’s trustworthiness.
@@
Structure. The structure of this paper is organized as follows. Section 2 reviews related work, discussing previous research on AI trustworthiness and classifying them. Section 3 provides the background, detailing subjective logic, its operators, and neural network calibration. Section 4 outlines the problem statement, highlighting the challenges in quantifying AI trustworthiness, especially in the context of neural networks. Section 5 describes our contribution, elaborating on the proposed method, the process of clustering predicted probabilities, and computing trustworthiness opinions. Finally, Section 6 presents the evaluation, including empirical analysis and results, demonstrating the effectiveness of our approach in real-world scenarios.
@@ 2 Related Work
The quantification of AI trustworthiness is an increasingly critical area of research, driven by the imperative to ensure that AI systems are reliable, fair, and interpretable. This section reviews seminal works that contribute to the development of methodologies and frameworks for evaluating and quantifying trust in AI systems, particularly focusing on deep neural networks (DNNs) and neural networks (NNs).
@@
Interpretable Trust Quantification Metrics for DNNs Wong et al. (2021): Wong, Wang, and Hryniowski propose a novel set of interpretable metrics aimed at quantifying the trustworthiness of DNNs based on their performance in a series of questions. Their research, outlined in "How Much Can We Really Trust You? Towards Simple Interpretable Trust Quantification Metrics for Deep Neural Networks," Wong et al. (2021) introduces metrics such as Question-Answer Trust, Trust Density, Trust Spectrum, and NetTrustScore. These metrics focus on the accuracy and confidence of AI outputs, emphasizing the importance of output confidence in the broader context of AI reliability without delving into other aspects like privacy or security. Their approach is static, as it establishes trust in the AI system through a one-time evaluation. It’s a black box method, focusing on the AI system’s outputs without requiring insight into its internal processes.
@@
Comprehensive Framework for Trustworthiness Evaluation Celdran et al. (2023): Contrasting with the singular focus on output confidence, Huertas Celdran and colleagues in "A Framework Quantifying Trustworthiness of Supervised Machine and Deep Learning Models," Celdran et al. (2023) offer a broader evaluation encompassing fairness, explainability, robustness, and accountability. This framework provides a structured methodology for a holistic measure of trust in AI systems, underlining the criticality of these four pillars in establishing a comprehensive understanding of AI trustworthiness. Their approach is static and does not explicitly align with the white or black box paradigms but offers a structured evaluation that could incorporate elements of both.
@@
Survey on Uncertainty in DNNs Gawlikowski et al. (2022): Gawlikowski et al. in "A Survey of Uncertainty in Deep Neural Networks," Gawlikowski et al. (2022) focus on the quantification of uncertainty inherent in DNN predictions. By categorizing uncertainty into model, data, and distributional types, their work highlights the importance of distinguishing these uncertainties to enhance the reliability and trustworthiness of AI systems in high-stakes applications. The survey underscores the challenge of establishing trust based solely on output accuracy/confidence. The paper delves into "Single Deterministic Methods," exploring a dynamic and white box approach for assessing uncertainty in neural network (NN) predictions. This method involves utilizing single deterministic network methods to quantify uncertainty through a single forward pass within the network, enabling dynamic evaluation of uncertainty in each prediction. Emphasizing confidence in output accuracy over other aspects of AI trustworthiness like privacy or security, it adopts internal uncertainty quantification approaches to emulate Bayesian modeling for quantifying model uncertainty. By directly assessing uncertainty from the network’s output, this methodology enhances interpretability and reliability of NN outputs, contributing significantly to discussions on AI trustworthiness, particularly in contexts where accurate and confident predictions are crucial.
@@
Evidential Deep Learning for Uncertainty Quantification Sensoy et al. (2018): Sensoy, Kaplan, and Kandemir introduce a novel framework in "Evidential Deep Learning to Quantify Classification Uncertainty" Sensoy et al. (2018) that leverages subjective logic for direct modeling of uncertainty in NN predictions. By treating predictions as subjective opinions and using Dirichlet distribution for uncertainty quantification, their method provides a dynamic and interpretable approach that improves robustness against adversarial perturbations and out-of-distribution queries. As previous work, this one is also a white-box approach and focusing only on uncertainty (accuracy/confidence). The main difference compared to previous work is that this work is based on Subjective Logic framework.
@@
DeepTrust Framework for Trust Quantification in NNs Cheng et al. (2020): Cheng, Nazarian, and Bogdan propose DeepTrust, a framework detailed in "There Is Hope After All: Quantifying Opinion and Trustworthiness in Neural Networks," which uses Subjective Logic to quantify trust in neural networks. DeepTrust is notable for its dynamic evaluation of trust, both in individual predictions and the neural network as a whole. By incorporating the trustworthiness of the dataset and the algorithmic processes, DeepTrust offers a nuanced assessment of AI trustworthiness. In a nutshell the solution proposed is a white box approach and is dynamic since it allows to evaluate trust in each prediction of the NN but still, it also allows to evaluate only trust in the NN. It can be applied to different property of trust such as accuracy, bias… However authors did not mention any property of AI Trustworthiness.
Method Black/White Box Dynamic/Static Subjective Logic Property
Wong et al. (2021) Black Static No Accuracy
Celdran et al. (2023) both Static No Trust
Single Deterministic Method in Gawlikowski et al. (2022) White Dynamic No Accuracy
Sensoy et al. (2018) White Dynamic Yes Accuracy
Cheng et al. (2020) White Dynamic Yes Trust
Our Approach Black Static Yes Accuracy
Table 1: Summary of Related Work (Transposed)
@@
Unlike some of the methods reviewed, our approach is static, as it evaluates the trustworthiness of the neural network as a whole rather than dynamically assessing the trustworthiness of individual predictions or uncertainties. By adopting a black-box approach, our method does not require insight into the internal workings of the neural network, making it broadly applicable to various types of models. Furthermore, our primary focus is on improving the accuracy of predicted probabilities through calibration, specifically addressing the alignment between model confidence and actual outcomes. While subjective logic is used to provide an intuitive and interpretable trustworthiness assessment, the emphasis remains on ensuring that the overall confidence of the network is reliable for deployment in critical applications.
3 Background
3.1 Subjective Logic
3.1.1 Subjective Opinion
@@
Subjective logic Jø sang (2016) is an extension of probabilistic logic to consider uncertainty. By adding uncertainty, subjective logic allows random variable to take a value such as "I don’t know". Subjective opinion notation is $\omega_{X}^{A}$, the subscript $X$ indicates the target variable or proposition to which the opinion applies, and the superscript $A$ indicates the subject agent who holds the opinion, i.e., the belief owner. In Subjective Logic we can distinguish three types of opinions (we consider that the domain of $X$ is $\mathbb{X}$):
  1. 1.
    binomial:
    • $card(\mathbb{X})=2(\mathbb{X}=\{x,\bar{x}\})$, $\omega_{X}^{A}=(b_{x},d_{x},u_{x},a_{x})$,
    • $b_{x}$ belief mass in support of $X$ being TRUE (i.e. $X=x$ ),
    • $d_{x}$ disbelief mass in support of $X$ being FALSE (i.e. $X=\bar{x}$ ),
    • $u_{x}$, uncertainty mass representing the vacuity of evidence,
    • $a_{x}$: base rate, i.e. prior probability of $x$ without any evidence,
    • $b_{x}+d_{x}+u_{x}=1$, Projected probability: $P(x)=b_{x}+a_{x}u_{x}$.
  2. 2.
    multinomial:
    • $card(\mathbb{X})>2$, $\omega_{X}^{A}=(\mathbf{b_{X}},u_{X},\mathbf{a_{X}})$,
    • $\mathbf{b_{X}}$ belief mass distribution over $\mathbb{X}$ ,
    • $u_{X}$, uncertainty mass representing the vacuity of evidence,
    • $\mathbf{a_{X}}$: base rate distribution over $\mathbb{X}$,
    • $\sum_{x}b_{x}+u_{X}=1$, Projected probability: $P(x)=b_{X}(x)+a_{X}(x)u_{X}$.
  3. 3.
    hypernomial:
    • $\mathbb{X}$ the domain and $\mathcal{R}(\mathbb{X})$ the corresponding hyperdomain. $card(\mathbb{X})>2$, $card(\mathcal{R}(\mathbb{X}))=2^{card(\mathbb{X})}$, $\omega_{X}^{A}=(\mathbf{b_{X}},u_{X},\mathbf{a_{X}})$,
    • $\mathbf{b_{X}}$ belief mass distribution over $\mathcal{R}(\mathbb{X})$ ,
    • $u_{X}$, uncertainty mass representing the vacuity of evidence,
    • $\mathbf{a_{X}}$: base rate distribution over $\mathcal{R}(\mathbb{X})$,
    • $\sum_{x}b_{x}+u_{X}=1$, Projected probability: $P(x)=b_{X}(x)+a_{X}(x)u_{X}$.
3.1.2 Subjective Logic Fusion Operators
@@
Subjective Logic can be used to analyze trust. Josang defined trust as a subjective binomial opinion $(t,d,u)$: (trust, distrust, untrust). He also introduced fusion operators that can be used to fuse information derived using different ways. These fusion operators are essential for merging opinions in various scenarios, such as trust analysis, decision-making, and expert systems.
@@
3.1.3 Binomial Opinion from evidence
In subjective logic, a binomial opinion is a probabilistic framework that expresses belief, disbelief, and uncertainty regarding a binary event or proposition. This opinion is particularly useful for decision-making under uncertainty, as it models not only the likelihood of an event but also the uncertainty due to limited or unreliable evidence.
The binomial opinion is characterized by four parameters:
  • $b_{x}$: The belief mass in support of proposition $X=x$ being true.
  • $d_{x}$: The disbelief mass, indicating the belief that $X=x$ is false.
  • $u_{x}$: The uncertainty mass, representing the vacuity of evidence for $X$.
  • $a_{x}$: The base rate, which is a prior probability expressing an agent’s bias toward $X$ in the absence of evidence.
These parameters follow the constraint $b_{x}+d_{x}+u_{x}=1$, ensuring the total belief, disbelief, and uncertainty add up to unity. The projected probability of $X$ is then defined as:
$P(x)=b_{x}+u_{x}\cdot a_{x}$
Equation 1 establishes a bijective mapping between a binomial opinion and the Beta probability density function (Probability of probability). The Beta distribution is a conjugate prior for binomial distributions in Bayesian statistics, and it models the probability of a binomial outcome based on observed evidence. The parameters of the Beta PDF are linked to the parameters of the binomial opinion through this mapping.
For a binomial opinion $\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$, the corresponding Beta Probability Density Function (PDF), denoted as $\text{Beta}^{e}(p_{x},r_{x},s_{x},a_{x})$, is described by two shape parameters $r_{x}$ and $s_{x}$, and a base rate $a_{x}$. The shape parameters $r_{x}$ and $s_{x}$ control the shape of the Beta distribution, while $a_{x}$ represents the prior belief about the binomial event in the absence of data.
The bijective mapping is as follows:
where $W$ is a weight parameter that controls the amount of vacuous evidence. In the case of vacuous opinions where no prior evidence exists, the non-informative default weight $W=2$ is commonly used, which corresponds to a uniform Beta distribution, $\text{Beta}(p_{x},1,1)$.
This mapping is important for translating subjective logic-based binomial opinions into Beta PDFs, which can then be used in Bayesian inference. For instance, the Beta PDF derived from a binomial opinion can serve as a prior distribution in cases where the opinion is used to infer future events based on past observations. This is especially relevant in scenarios such as decision-making systems, neural networks, and probabilistic forecasting models, where belief updating based on new evidence is crucial.
3.2 Neural Network Calibration
Neural networks have become the backbone of many modern AI applications due to their ability to learn and generalize from large datasets. However, an important aspect of their performance that has garnered attention is calibration. Calibration refers to the degree to which a model’s predicted probabilities align with the true likelihood of outcomes. A well-calibrated model provides predicted probabilities that match the actual frequencies of events. For example, among all instances where a model predicts a 70% probability of an event, that event should occur approximately 70% of the time.
Despite their high accuracy, many state-of-the-art neural networks, such as deep convolutional networks used in image classification or recurrent networks in natural language processing, often suffer from poor calibration. This misalignment can lead to overconfident or underconfident predictions, which is problematic in applications where uncertainty estimation is crucial, such as medical diagnosis or autonomous driving.
3.2.1 Calibration Error
@@
While useful, ECE does not offer a full view of a model’s trustworthiness because it focuses solely on probability alignment without considering the underlying uncertainty in predictions.
3.2.2 Calibration Techniques
Several methods have been developed to improve neural network calibration. One of the most effective and widely-used techniques is Temperature Scaling, a simple post-processing method introduced by Guo et al. Guo et al. (2017). In this method, the logits (the raw outputs of the neural network before applying the softmax function) are divided by a scalar temperature parameter $T$, which is optimized using a validation set. The modified logits are then fed into the softmax function to generate calibrated probabilities. The temperature parameter $T$ is tuned to reduce the difference between predicted probabilities and actual outcomes, as measured by metrics like the Negative Log-Likelihood (NLL).

### v2

# Quantifying Calibration Error in Neural Networks through Evidence-Based Theory thanks: We thank the partners in the European CONNECT project for their fruitful discussions on trust management. This work is partially funded by the HORIZON CONNECT project under EU grant agreement no. 101069688

@@ Abstract
Trustworthiness in neural networks is crucial for their deployment in critical applications, where reliability, confidence, and uncertainty play a pivotal role in decision-making. Traditional performance metrics such as accuracy and precision fail to capture these aspects, particularly in cases where models exhibit overconfidence. To address these limitations, this paper introduces a novel framework for quantifying the trustworthiness of neural networks by incorporating subjective logic into the evaluation of Expected Calibration Error (ECE). This method provides a comprehensive measure of trust, distrust, and uncertainty by clustering predicted probabilities and fusing opinions using appropriate fusion operators. We demonstrate the effectiveness of this approach through experiments on the MNIST and CIFAR-10 datasets, where post-calibration results indicate improved trustworthiness. The proposed framework offers a more interpretable and nuanced assessment of AI models, with potential applications in sensitive domains such as healthcare and autonomous systems.
Index Terms:
Neural Network Trustworthiness, Subjective Logic, Confidence Calibration, Expected Calibration Error (ECE), Belief Fusion
I Introduction
Artificial Intelligence (AI) systems, particularly neural networks, are increasingly employed in critical applications such as healthcare, finance, and autonomous systems. These systems play an integral role in decision-making, where the trustworthiness of their predictions becomes paramount. Trustworthiness in AI encompasses attributes such as reliability, robustness, fairness, and transparency, yet these qualities are often difficult to evaluate, particularly in neural networks, which are typically viewed as ”black-box” models. This opacity raises significant concerns about their trustworthiness, especially in sensitive domains where incorrect decisions can lead to severe consequences.
@@ I Introduction
Problem. Evaluating the trustworthiness of an AI system, particularly neural networks, poses significant challenges. Traditional metrics like accuracy, precision, and recall do not capture the uncertainties and subjective aspects of trustworthiness. Confidence calibration methods, such as those discussed in [5], introduce the Expected Calibration Error (ECE) to align predicted probabilities with actual outcomes, aiming to make AI systems more trustworthy. However, ECE, while useful, is not easily interpretable and does not provide a comprehensive view of an AI system’s trustworthiness.
@@
Structure. This paper is organized as follows. Section 2 reviews related work, discussing previous research on AI trustworthiness and classifying them. Section 3 provides a background, detailing subjective logic, its operators, and neural network calibration. Section 4 outlines the problem statement, highlighting the challenges in quantifying AI trustworthiness, especially in the context of neural networks. Section 5 describes our contribution, elaborating on the proposed method, the process of clustering predicted probabilities, and computing trustworthiness opinions. Finally, Section 6 presents the evaluation, including empirical analysis and results, demonstrating the effectiveness of our approach in real-world scenarios.
@@ II Related Work
The quantification of AI trustworthiness is an increasingly critical area of research, driven by the imperative to ensure that AI systems are reliable, fair, and interpretable. This section reviews key works on evaluating trust in neural networks (NNs).
@@
Interpretable Trust Quantification Metrics for NNs [12]: proposes a novel set of interpretable metrics aimed at quantifying the trustworthiness of NNs based on their performance in a series of questions. Their research introduces metrics such as Question-Answer Trust, Trust Density, Trust Spectrum, and NetTrustScore. These metrics focus on the accuracy and confidence of AI outputs, emphasizing the importance of output confidence in the broader context of AI reliability without delving into other aspects like privacy or security. Their approach is static, as it establishes trust in the AI system through a one-time evaluation. It is a black box method, focusing on the AI system’s outputs without requiring insight into its internal processes.
@@
Comprehensive Framework for Trustworthiness Evaluation [1]: Contrasting with the singular focus on output confidence, [1] offers a broader evaluation encompassing fairness, explainability, robustness, and accountability. This framework provides a structured methodology for a holistic assessment of trust in AI systems, underlining the importance of these four pillars in establishing a comprehensive understanding of AI trustworthiness. Their approach is static and does not explicitly align with the white or black box paradigms but offers a structured evaluation that could incorporate elements of both.
@@
Survey on Uncertainty in NNs [4]: focuses on the quantification of uncertainty inherent in NN predictions. By categorizing uncertainty into model, data, and distributional types, their work highlights the importance of distinguishing these uncertainties to enhance the reliability and trustworthiness of AI systems in high-stakes applications. The survey underscores the challenge of establishing trust based solely on output accuracy/confidence. The paper delves into ”Single Deterministic Methods,” exploring a dynamic and white box approach for assessing uncertainty in neural network (NN) predictions. This method involves utilizing single deterministic network methods to quantify uncertainty through a single forward pass within the network, enabling dynamic evaluation of uncertainty in each prediction. Emphasizing confidence in output accuracy over other aspects of AI trustworthiness like privacy or security, it adopts internal uncertainty quantification approaches to emulate Bayesian modeling for quantifying model uncertainty. By directly assessing uncertainty from the network’s output, this methodology enhances interpretability and reliability of NN outputs, contributing significantly to discussions on AI trustworthiness, particularly in contexts where accurate and confident predictions are crucial.
@@
Evidential Deep Learning for Uncertainty Quantification [10]: introduces a novel framework that leverages subjective logic for direct modeling of uncertainty in NN predictions. By treating predictions as subjective opinions and using Dirichlet distribution for uncertainty quantification, their method provides a dynamic and interpretable approach that improves robustness against adversarial perturbations and out-of-distribution queries. Like previous work, this one is also a white-box approach and focusing only on uncertainty (accuracy/confidence). The main difference compared to previous work is that this work is based on Subjective Logic framework.
@@
DeepTrust Framework for Trust Quantification in NNs [2]: proposes DeepTrust which uses Subjective Logic to quantify trust in neural networks. DeepTrust is notable for its dynamic evaluation of trust, both in individual predictions and the neural network as a whole. By incorporating the trustworthiness of the dataset and the algorithmic processes, DeepTrust offers a nuanced assessment of AI trustworthiness. In a nutshell the solution proposed is a white box approach and is dynamic since it allows to evaluate trust in each prediction of the NN but still, it also allows to evaluate only trust in the NN.
@@
Unlike some of the methods reviewed, our approach is static, as it evaluates the trustworthiness of the neural network as a whole rather than dynamically assessing the trustworthiness of individual predictions or uncertainties. By adopting a black-box approach (meaning that the AI to assess is seen as a black box), our method does not require insight into the internal workings of the neural network, making it broadly applicable to various types of models. Furthermore, our primary focus is on improving the accuracy of predicted probabilities through calibration, specifically addressing the alignment between model confidence and actual outcomes. While subjective logic is used to provide an intuitive and interpretable trustworthiness assessment, the emphasis remains on ensuring that the overall confidence of the network is reliable for deployment in critical applications.
III Background
III-A Subjective Logic
III-A1 Subjective Opinion
@@
Subjective logic [6] is an extension of probabilistic logic that incorporates a second level of uncertainty. This additional layer allows for expressions such as ”I don’t know,” enabling a random variable to capture uncertainty more explicitly. Moreover, subjective logic allows an agent to express an opinion about a given proposition. The notation for a subjective opinion is $\omega_{X}^{A}$. The subscript $X\in\mathbb{X}$ indicates the target variable or proposition to which the opinion applies, and the superscript $A$ indicates the agent (i.e., the belief owner) who holds the opinion. In Subjective Logic, depending on $\mathbb{X}$ we can distinguish three types of opinions: binomial, multinomial and hypernomial. For our work, only the binomial type is relevant.
In binomial opinion, the cardinal of the domain $\mathbb{X}$ is 2 ($\left(\{x,\bar{x}\}\right)$ and the opinion has the form $\omega_{X}^{A}=(b_{x},d_{x},u_{x},a_{x})$, where $b_{x}$ is the belief mass in support of $x$ being TRUE (i.e. $X=x$ ), $d_{x}$ is the disbelief mass in support of $X$ being FALSE (i.e. $X=\bar{x}$ ), $u_{x}$ is the uncertainty mass representing the vacuity of evidence, and $a_{x}$ is the base rate, i.e. prior probability of $x$ without any evidence. These parameters follow the constraint $b_{x}+d_{x}+u_{x}=1$, ensuring the total belief, disbelief, and uncertainty add up to unity.
The projected probability of the opinion is then defined as:
$P(x)=b_{x}+u_{x}\cdot a_{x}$
Equation 1 establishes a bijective mapping between a binomial opinion and the Beta probability density function (Probability of probability). This Beta Probability Density Function, denoted as $\text{Beta}^{e}(p_{x},r_{x},s_{x},a_{x})$, is described by two shape parameters $r_{x}$ and $s_{x}$, and a base rate $a_{x}$. The shape parameters $r_{x}$ and $s_{x}$ control the shape of the Beta distribution, while $a_{x}$ represents the prior belief about the binomial event in the absence of data. $r_{x}$ and $s_{x}$ can be interpreted as, respectively, the number of positive and negative evidence for $X$ taking the value $x$.
The bijective mapping is as follows [6]:
where $W$ is a weight parameter that controls the amount of vacuous evidence. In the case of vacuous opinions where no prior evidence exists, the non-informative default weight $W=2$ is commonly used, which corresponds to a uniform Beta distribution, $\text{Beta}(p_{x},1,1)$. More binomial opinion quantification method can be found in another of our submitted paper [8].
III-A2 Analyze Trust with Subjective Logic
@@
Subjective logic can be used to analyze trust [7, 9]. [6] defines trust as a subjective binomial opinion $(t,d,u)$ [^1] The base rate is usually always set to $0.5$., where $t$ is the belief mass (or trust mass), $d$ is the disbelief mass (or distrust mass), and $u$ is the uncertainty (or untrust mass). It also introduced fusion operators that can be used to fuse information derived using different ways. These fusion operators are essential for merging opinions in various scenarios, such as trust analysis and decision-making.
@@
III-B Neural Network Calibration
Neural networks have become the backbone of many modern AI applications due to their ability to learn and generalize from large datasets. However, an important aspect of their performance that has received increasing attention is calibration. Calibration refers to the degree to which a model’s predicted probabilities align with the true likelihood of outcomes. A well-calibrated model provides predicted probabilities that match the actual frequencies of events. For example, among all instances where a model predicts a 70% probability of an event, that event should occur approximately 70% of the time.
Despite their high accuracy, many state-of-the-art neural networks, such as deep convolutional networks used in image classification or recurrent networks in natural language processing, often suffer from poor calibration [11]. This misalignment can lead to overconfident or underconfident predictions, which is problematic in applications where uncertainty estimation is crucial, such as medical diagnosis or autonomous driving.
III-B1 Calibration Error
@@
While useful, ECE does not offer a full view and interpretable quantity of a model’s trustworthiness.
III-B2 Calibration Techniques
Several methods have been developed to improve neural network calibration. One of the most effective and widely-used techniques is Temperature Scaling, a simple post-processing method introduced by [5]. In this method, the logits [^2] The logits are the raw outputs of the neural network before applying the softmax function are divided by a scalar temperature parameter $T$, which is optimized using a validation set. The modified logits are then fed into the softmax function to generate calibrated probabilities. The temperature parameter $T$ is tuned to reduce the difference between predicted probabilities and actual outcomes, as measured by metrics like the Negative Log-Likelihood (NLL).